Sir,i ll give you the scenario,One question,4 options...he can click any options until the answer turns to true...ie it can be FFFT or FFTT etc....I want to know how accurate he is in solving this problem....please help me...
Answer by sudhanshu_kmr(1152)


total number of ways = 2^4 = 16
probability of right sequence = 1/16